OpenStudy (anonymous):

what is 10/4-54/4 equal?

OpenStudy (anonymous):

-11 10/4-54/4= -44/4=-11

sammixboo (sammixboo):

\(\large\color{magenta}{\frac{10}{4} - \frac{54}{4}}\) = \(\large\color{magenta}{\frac{-44}{1}}\)

sammixboo (sammixboo):

Because \(\sf\large\color{magenta}{10 - 54 = -41,~and~the~4~should~stay~the~same.}\)

sammixboo (sammixboo):

And yes I did, lol.

sammixboo (sammixboo):

Omg sooo sorry -o-

sammixboo (sammixboo):

I am like half awake atm sooooo

sammixboo (sammixboo):

I am just trying to answer 7 more questions to reach 700 questions answered -,-

sammixboo (sammixboo):

Yep

OpenStudy (anonymous):

Ooh, so long ago T-T